OpenStudy (ledah):

Solve using substitution 6x + 2y = 0 4x + y = -1

OpenStudy (anonymous):

start by solving the second equation \(4x+y=-1\) for \(y\) to get \(y=-1-4x\) then substitute in to first equation to get \[6x+2(-1-4x)=0\] and solve that one

OpenStudy (anonymous):

\[6x-2-8x=0\]\[-2x-2=0\]\[2x=-2\] \[x=-1\] etc

OpenStudy (anonymous):

thankyou so much !